### v1

# Incremental Learning for Robot Shared Autonomy

Abstract
Shared autonomy holds promise for improving the usability and accessibility of assistive robotic arms, but current methods often rely on costly expert demonstrations and lack the ability to adapt post-deployment. This paper introduces ILSA, an Incrementally Learned Shared Autonomy framework that continually improves its assistive control policy through repeated user interactions. ILSA leverages synthetic kinematic trajectories for initial pretraining, reducing the need for expert demonstrations, and then incrementally finetunes its policy after each manipulation interaction, with mechanisms to balance new knowledge acquisition with existing knowledge retention during incremental learning. We validate ILSA for complex long-horizon tasks through a comprehensive ablation study and a user study with 20 participants, demonstrating its effectiveness and robustness in both quantitative performance and user-reported qualitative metrics. Code and videos are available at https://ilsa-robo.github.io/.
I INTRODUCTION
Assistive devices such as wheelchair-mounted robotic arms have the potential to benefit millions of individuals with physical disabilities [36] and enhance their independence [5, 2, 27]. However, controlling assistive robotic arms via the current standard of teleoperation is challenging, especially for those with limited mobility. Shared autonomy, where human operators and autonomous systems share decision-making, can make these devices more intuitive and effective for complex tasks, enhancing performance and safety [33, 13].
Recent shared autonomy systems have begun employing imitation learning, which learns an assistive control policy from human demonstrations [23, 20, 16, 8, 39]. However, training these models often requires substantial “expert” demonstrations, which are costly and difficult to scale [14, 29, 40]. Additionally, most shared autonomy and imitation learning implementations focus on creating robust systems before deployment but often remain static once in use, lacking mechanisms for continuous improvement. Without the ability to adapt post-deployment, these systems are limited by their initial performance.
(a) Cereal pouring task, trial 1
(b) Cereal pouring task, trial 4
(c) Pill bottle storage task, trial 1
(d) Pill bottle storage task, trial 4
Figure 1: We introduce an Incrementally Learned Shared Autonomy (ILSA) system that is initially pretrained on synthetic kinematic trajectories and then continuously refined as the user repeatedly interacts with it. Left: During the initial trials, ILSA lacks the knowledge required to avoid collisions, leading to task objects colliding with obstacles. Right: With only a few interactions, by the fourth trials, ILSA has learned effective collision avoidance from past user interactions, assisting the user in completing tasks more smoothly.
Incremental learning [37, 12] offers a promising solution to these challenges by enabling neural networks to continuously learn and improve from non-stationary streams of data. This paradigm is particularly well-suited for shared control systems, which are used repeatedly and generate continual interaction data.
Building on these insights, we introduce an Incrementally Learned Shared Autonomy (ILSA) framework (Fig. 2), which improves its generated assistive actions after each user interaction. ILSA begins with a pretraining phase using synthetic kinematic trajectories generated based on simplistic and intuitive rules as described in Section III-B, thereby eliminating the need for actual expert demonstrations. This pretraining phase enables ILSA to provide a basic level of assistance the first time a user uses it. After pretraining, ILSA is incrementally finetuned with each interaction. We design the finetuning process of ILSA to balance learning from new data and retaining existing knowledge. (Refer to Fig. 1 for an illustration of ILSA’s learning progression.) We demonstrate the effectiveness of the designs through a quantitative ablation study.
Figure 2: Our Incrementally Learned Shared Autonomy (ILSA) framework. $a_{m,t}$ is the intermediate output action without user guidance, generated by $\mathcal{F^{\prime}}_{\theta^{\prime}}(s_{t})$. $a_{f,t}$ is the final output action with user guidance, generated by ′′$\mathcal{F^{\prime\prime}}_{\theta^{\prime\prime}}(s_{t},a_{u,t})$. $\mathcal{F^{\prime}}_{\theta^{\prime}}$ and ′′$\mathcal{F^{\prime\prime}}_{\theta^{\prime\prime}}$ are subsets of the overall action generation model $\mathcal{F}_{\theta}(s_{t},a_{u,t})$, with blue and green annotations indicating their components, respectively. Their roles are detailed in Section III-C. $L_{demo,m}$, $L_{demo,f}$, $L_{order}$, and $L_{direc}$ are loss terms designed to balance task completion and the user’s sense of control, as defined in Section III-C. The red components highlight the training designs used during the incremental learning phase.
To formally evaluate ILSA, we conducted a user study where 20 participants used a Kinova robotic arm to perform two complex, long-horizon tasks. Study results support our hypothesis that (1) ILSA supports faster task completion and easier manipulator control as compared to pure teleoperation; and (2) ILSA improves task performance over time as a user repeatedly performs a task, in contrast to a static shared autonomy method.
Our contributions are summarized as follows:
  • We propose ILSA, a novel Incrementally Learned Shared Autonomy framework that improves a shared control policy through continual user interactions without extensive expert demonstrations.
  • Through a quantitative ablation study, we investigate how ILSA balances learning from new data while retaining existing knowledge during finetuning.
  • We conducted a human study, demonstrating that ILSA rapidly enhances task performance over time, significantly outperforming teleoperation and non-incremental shared autonomy.
II RELATED WORK
II-A Shared Autonomy
Despite varying definitions and settings, most shared autonomy approaches aim to map human inputs—often suboptimal due to limited knowledge or experience [31, 39], or constrained to a lower dimension than the robot’s action space [23, 15]—into robot movements that successfully accomplish the task. Regardless of the specific approach, existing research primarily focuses on developing a robust system before deployment, which often requires a complex expert demonstration collection process, especially in recent learning-based methods [14, 29, 20, 8, 39]. This focus likely stems from the perspective that human operators in shared autonomy systems are typically non-experts in need of assistance. However, we contend that even non-expert human operators can contribute positively to the task, and their intelligence should be harnessed. Thus, we propose ILSA, a framework that continuously evolves and improves through repeated user interaction.
II-B Human-in-the-loop Machine Learning
Another related area within full autonomy is human-in-the-loop machine learning [9, 1], where an autonomous agent learns from human interactions to improve performance. Various forms of human feedback, such as demonstrations, preferences, rankings [4], interventions [17, 25, 35], and scalar-valued critiques [24, 38], are integrated into these processes.
Among these works, interactive imitation learning [6, 28, 3], which incorporates human interventions into imitation learning, is particularly relevant to our work. However, interactive imitation learning methods aim to develop fully autonomous systems using expert interventions and still require extensive expert demonstrations to train on before deployment [17, 25, 34, 7, 22, 21]. In contrast, we focus on a shared autonomy system that assists non-expert operators, pretrained on synthetic kinematic trajectories and finetuned through interactions with users. Moreover, existing methods often need numerous policy deployments and substantial human feedback before each update, limiting practical use. Our method updates the policy after each use by the operator, balancing learning from new data while retaining existing knowledge.
II-C Incremental Learning
Incremental learning, sometimes referred to interchangeably with continual learning and lifelong learning, aims to enable data-driven models to incrementally learn from non-stationary streams of data [12]. A core challenge in incremental learning is mitigating a phenomenon known as catastrophic forgetting [26, 10], where new learning can cause previously acquired capabilities to be lost.
The field of incremental learning has seen significant growth in recent years, with applications in various areas such as medical diagnosis [19], autonomous driving [32], and financial forecasting [30]. However, there is minimal work that incorporates incremental learning into robot shared autonomy. Our approach seeks to bridge this gap by leveraging incremental learning to continuously enhance the shared autonomy system’s performance over time through ongoing interactions with human operators.
III Methodology
III-A Problem Statement
We formulate our shared autonomy task as a sequential decision process defined by tuple $(\mathcal{S},\mathcal{A}_{u},\mathcal{A}_{r},\mathcal{T})$, where $s_{t}\in\mathcal{S}$ denotes the task state at time step $t$, $a_{u,t}\in\mathcal{A}_{u}$ denotes the user action, $a_{r,t}\in\mathcal{A}_{r}$ denotes the robot action (we use the 6-DoF delta end-effector position, including Cartesian coordinates and Euler angles). $\mathcal{T}$: $\mathcal{S}\times\mathcal{A}_{r}\rightarrow\mathcal{S}$ is an unobserved transition function.
Our goal is to identify a function $\mathcal{F}_{\theta}(s_{t},a_{u,t})$: $\mathcal{S}\times\mathcal{A}_{u}\rightarrow\mathcal{A}_{r}$ that maps the task state and user action to a robot action that is both task-optimal and aligned with the user’s intent. While $\mathcal{F}_{\theta}$ may be imperfect initially, the objective is to improve it over time as the user interacts with the shared autonomy system.
III-B Synthetic Kinematic Trajectories Generation
We use synthetic kinematic trajectories to warm-start ILSA, reducing the need for expert demonstrations. To generate these trajectories, we divide multi-step tasks into distinct subtasks. For each subtask, we randomly sample task object locations, calculate the corresponding target end-effector positions, and use linear interpolation between initial and target end-effector positions to create trajectories.
For example, in a cereal-pouring task divided into three subtasks—(1) grasp the cereal container, (2) move the container next to the bowl, and (3) pour the cereal—we randomly sample locations for the cereal container and bowl within the robot’s workspace. Based on these object locations, we compute the robot end-effector’s target positions for each subtask. Sequences of end-effector positions are then generated via linear interpolation between initial and target end-effector positions, forming task states $s_{t}$. Differences between successive positions yield the robot actions $a_{r,t}$. Note that these kinematic trajectories may be imperfect and might not account for factors like obstacle avoidance. However, warm-starting ILSA with diverse trajectories supports broader generalization capabilities without needing to collect expert demonstrations for each new task.
These synthetic kinematic trajectories yield tuples $(s_{t},a_{r,t})$.
To train ILSA on kinematic trajectories, we also generate synthetic user actions $a_{u,t}$ to mimic real user inputs. We design $a_{u,t}$ as a three-dimensional vector, with each value representing the user’s intended end-effector movement in the $x$, $y$, and $z$ translational dimensions. To generate such $a_{u,t}$, for each dimension $i\in\{x,y,z\}$, if the target end-effector position in dimension $i$ is greater than the current position in the trajectory, we sample $a_{u,t,i}$ to be in $[0,1]$, otherwise we sample $a_{u,t,i}$ in $[-1,0]$. In our real-world experiments, the user provides $a_{u,t}$ using two joysticks on an Xbox controller. One joystick operates two axes of end effector motion (e.g., $x$- and $y$-axes in translation mode) and the other joystick operates the third axis.
In summary, we generate trajectories each consisting of a series of $(s_{t},a_{u,t},a_{r,t})$ triplets. We generate 50 trajectories for each task.
III-C Robot Action Generation Model
Our robot action generation model $\mathcal{F}_{\theta}(s_{t},a_{u,t})$, shown in Fig. 2, produces two types of actions at each time step: an intermediate action $a_{m,t}=\mathcal{F^{\prime}}_{\theta^{\prime}}(s_{t})$ without user guidance and a final action $a_{f,t}=\mathcal{F^{\prime\prime}}_{\theta^{\prime\prime}}(s_{t},a_{u,t})$ with user guidance. Here, $\mathcal{F^{\prime}}_{\theta^{\prime}}$ and ′′$\mathcal{F^{\prime\prime}}_{\theta^{\prime\prime}}$ represent different subsets of the overall model $\mathcal{F}_{\theta}$, as illustrated in Fig. 2. While only the action with user guidance is executed, having an intermediate output is crucial during the incremental learning phase, as we will discuss in Section III-D.
We pretrain $\mathcal{F}_{\theta}$ in an end-to-end fashion with synthetic kinematic trajectories (Section III-B). We achieve the balance between task completion and the user’s sense of control through a loss function defined as $\mathcal{L}=\alpha\mathcal{L}_{demo,m}+\beta\mathcal{L}_{demo,f}+\gamma\mathcal{L}_{direc}+\delta\mathcal{L}_{order}$, where $\alpha$, $\beta$, $\gamma$ and $\delta$ are weights that adjust the contribution of each loss term. We set $\alpha=\beta=1$ and $\gamma=\delta=100$.
$\mathcal{L}_{demo,m}$ and $\mathcal{L}_{demo,f}$ are defined as follows:
$\mathcal{L}_{demo,m}=\|a_{m,t}-a_{r,t}\|_{2}^{2}$ (1)
$\mathcal{L}_{demo,f}=\|a_{f,t}-a_{r,t}\|_{2}^{2}$ (2)
where $a_{r,t}$ is the robot action in the kinematic trajectories. These two loss terms supervise the intermediate and final outputs respectively, guiding the robot to take actions towards task completion. To further ensure that the final executed action aligns with the user’s input, we introduce two additional loss terms, action direction loss $\mathcal{L}_{direc}$ and action ordering loss $\mathcal{L}_{order}$, to supervise $a_{f,t}$.
$\mathcal{L}_{direc}$ ensures that the signs of the robot action align with the user action in each dimension. It is defined as:
$\mathcal{L}_{direc}=\sum_{i\in\{x,y,z\}}|a_{f,t,i}|\cdot 1_{\mathbb{R}_{<0}}(a_{f,t,i}\cdot a_{u,t,i})$ (3)
where $a_{f,t,i}$ and $a_{u,t,i},i\in\{x,y,z\}$ denote the final output action and the user action in the $x$, $y$, $z$ dimensions, respectively. $1_{\mathbb{R}_{<0}}()$ is the indicator function for the set of negative real numbers.
The action ordering loss, $\mathcal{L}_{order}$, ensures that the relative magnitudes of the robot action align with the relative magnitudes of the user action across the $x$, $y$, and $z$ dimensions. It is defined as:
$\mathcal{L}_{order}=\sum_{(i,j)\in P}$ $\ ||a_{f,t,i}|-|a_{f,t,j}||\cdot$ (4)
$1_{\mathbb{R}_{<0}}\Big{(}(|a_{f,t,i}|-|a_{f,t,j}|)(|a_{u,t,i}|-|a_{u,t,j}|)\Big{)}$
where $P=\{(x,y),(x,z),(y,z)\}$ represents the pairs of dimensions over which the relative magnitude differences are compared.
This term penalizes cases where the ordering of action magnitudes between the robot and user is inconsistent. For example, if the user’s action along the $x$-dimension is greater than along the $y$-dimension, but the robot’s corresponding action does not reflect this relationship, a penalty is applied. By minimizing this loss, the model is encouraged to maintain consistency with user’s input.
III-D Incremental Learning
ILSA starts with the robot action generation model $\mathcal{F}_{\theta}$ pretrained on kinematic trajectories. Our goal is to incrementally finetune $\mathcal{F}_{\theta}$ using interaction data collected during deployment.
During deployment, at each time step, we obtain the user action $a_{u,t}$ from the control interface and input it into $\mathcal{F}_{\theta}$ to generate the final robot action $a_{f,t}$. Since the pretrained model may produce errors, we compute the cosine similarity between $a_{u,t}$ and $a_{f,t}$. If the similarity is above an empirically set threshold (0.5), we execute $a_{f,t}$. If it falls below the threshold, indicating a discrepancy, we alert the user by briefly pausing the robot’s motion. If the user persists with the same input, the robot will execute $a_{u,t}$ instead of $a_{f,t}$.
As the robot continues to operate using this process, we collect interaction data to finetune $\mathcal{F}_{\theta}$. However, simply updating the entire model with these raw trajectories has limitations. In the following sections, we discuss these challenges and propose corresponding solutions.
III-D1 Corrected Trajectory Generation
Using raw trajectories of user interactions to update $\mathcal{F}_{\theta}$ can be suboptimal, as our experiments show that when users overwrite $a_{f,t}$ with $a_{u,t}$, it is typically to correct errors previously made by the system.
For example, in the case shown in Fig. 1(a), when the robot arm collides with a shelf while retrieving a cup containing cereal, the user wants to command the robot to move down to avoid further mistakes. This action is an outlier unexpected by the robot, hence the user needs to overwrite $a_{f,t}$ with $a_{u,t}$. This process is only necessary because previously, the pretrained model didn’t account for environmental collisions and mistakenly attempted a linear trajectory to move next to the bowl, forcing the user to correct it.
As illustrated in this example, raw trajectories contain both policy errors and user corrections that wouldn’t be necessary without the errors. Directly using these trajectories for training risks misleading the model into treating these corrective actions as optimal behavior to be replicated.
To address this, we generate a “corrected trajectory” for each raw trajectory observed during human use. As shown in Fig. 3, when the user overwrites $a_{f,t}$ (red segments in Fig. 3), let $t_{1}$ be the start of the overwrite and $t_{2}$ the time when the robot resumes executing $a_{f,t}$. Before $t_{1}$, if the robot executes $a_{f,t}$ for a period (blue segments), let $t_{0}$ be the start of that period. Let $p_{t_{0}}$, $p_{t_{1}}$, and $p_{t_{2}}$ represent the robot’s positions at these times.
Our assumption is that by correcting the robot position from $p_{t_{1}}$ to $p_{t_{2}}$, the user essentially believes the ideal trajectory goes directly from $p_{t_{0}}$ to $p_{t_{2}}$. Therefore, we simply take the linear trajectory between $p_{t_{0}}$ and $p_{t_{2}}$ (green segments) as the corrected trajectory segment. We then perform linear interpolation to generate new $(s_{t},a_{r,t})$ pairs and apply the same method from Section III-B to generate $a_{u,t}$, forming new $(s_{t},a_{u,t},a_{r,t})$ triplets for training.
Figure 3: A corrected trajectory (green segments) constructed from ILSA actions (blue segments) and user corrections (red segments) for the first trial of a task.
III-D2 Layered Supervision
After generating the corrected trajectory, a straightforward approach is to update the model using only the newly generated trajectory. However, for rapid learning, ILSA is designed to be updated after each new interaction, and training on only one or a few new trajectories can result in overfitting, gradient noise, and catastrophic forgetting [11, 26, 18].
Incorporating synthetic kinematic trajectories into the training set can help address this issue. Instead of directly combining the kinematic trajectories with new trajectories, which may lead to data imbalance, we propose a layered supervision approach:
  1. 1.
    Final Output Supervision: We only use the new user-generated trajectories to supervise $a_{f,t}$, i.e., finetune ′′$\mathcal{F^{\prime\prime}}_{\theta^{\prime\prime}}$.
  2. 2.
    Intermediate Output Supervision: Both the synthetic kinematics trajectories and new user-generated trajectories are used to supervise $a_{m,t}$, i.e., finetune $\mathcal{F^{\prime}}_{\theta^{\prime}}$.
This approach is illustrated by the red arrows in Fig. 2. It integrates the benefits of both synthetic kinematic trajectories and new trajectories, while prioritizing the model’s adaptation to the newly generated user data during fine-tuning. Compared to other commonly used techniques in human-in-the loop machine learning methods such as assigning different weights to different types of data [25, 22, 21], ILSA avoids the need for data reweighting or dataset subsampling.
III-D3 Partial Model Update
To improve the robustness of the finetuning phase, we update only the transformer in the robot action generation model, which maps the task state $s_{t}$ into state features, while freezing other parts of the architecture. The rationale is that the initial model’s inefficiencies mainly stem from the simplistic linear nature of the synthetic kinematic trajectories, which fail to take full consideration of all environmental factors. Updating the transformer, responsible for modeling environmental variations, can sufficiently address this limitation.
IV Experiments
IV-A Tasks and Experiment Settings
All experiments are conducted with a 6-DOF Kinova Gen3 robotic arm, and users use an Xbox controller to send commands to the robot. We test the effectiveness of our model on two complex, multi-stage tasks:
IV-A1 Cereal Pouring
Pick up a cup of cereal placed under a shelf and pour the cereal into a bowl on top of the shelf.
IV-A2 Pill Bottle Storage
Open a drawer, pick up a pill bottle behind the drawer, and place it back inside the opened drawer.
For each task, a complete experiment consists of four interactions, where the user completes the task four times with different object positions each time. We refer to these interactions as “trial 1”, “trial 2”, “trial 3”, and “trial 4”. ILSA updates $\mathcal{F}_{\theta}$ after each trial.
Figure 4: Average task completion times from our ablation study on the cereal pouring task. Error bars show standard deviations.
IV-B Ablation Study
Prior to the user study, we conducted an ablation study on the cereal pouring task to verify the design choices of ILSA. The ablated components and alternative designs evaluated are:
  • ILSA finetunes on corrected trajectories: A. Finetune on original uncorrected trajectories.
  • ILSA uses new data to supervise both $\bm{a_{f,t}}$ and $\bm{a_{m,t}}$, and kinematic trajectories to supervise only $\bm{a_{m,t}}$: B1. Use new data only to supervise both $a_{f,t}$ and $a_{m,t}$; B2. Use both new data and kinematic trajectories to supervise both $a_{f,t}$ and $a_{m,t}$; B3. Use new data and randomly selected 10 kinematic trajectories to supervise both $a_{f,t}$ and $a_{m,t}$; B4. Use both new data and kinematic trajectories to supervise both $a_{f,t}$ and $a_{m,t}$, reweighted with $w_{new}=0.5\cdot\frac{|D_{new}|+|D_{kinematic}|}{|D_{new}|}$ and $w_{kinematic}=0.5\cdot\frac{|D_{new}|+|D_{kinematic}|}{|D_{kinematic}|}$, where $|D_{new}|$ and $|D_{kinematic}|$ represent the number of data points in the new and kinematic trajectories, respectively (This design is inspired by human-in-the-loop machine learning works [25, 22]).
  • ILSA updates only the Transformer: C1. Update the entire architecture; C2. Freeze the transformer and update other networks.
For the ablation study, one researcher runs five experiments for each model on the cereal pouring task. Each experiment consists of four trials with different task layouts. The task completion times of different models are shown in Fig. 4.
As shown in Fig. 4, our proposed method, ILSA, consistently decreases task completion times across trials. With just a few user interactions, ILSA lowers the average task completion time from 55.2 seconds in trial 1 to 35.6 seconds in trial 4, which is a 35.5% reduction. From trial 2 to trial 4, ILSA has the lowest average task completion time and one of the smallest standard deviations, demonstrating both effectiveness and repeatability.
In contrast, the method using uncorrected trajectories (A1) shows a smaller reduction in task completion time in trials 2 and 3, with larger standard deviations. This is mainly because user corrections for previous errors in the uncorrected trajectories are outliers that can cause overfitting. This can lead to errors in later trials, resulting in higher average task completion times and standard deviations.
Similarly, updating the model only on new data (B1), updating all neural networks in the architecture (C1), or freezing the transformer and updating other networks (C2) also leads to worse incremental learning and task performance. These ablated methods performed poorly on task layouts that are highly dissimilar to previously seen scenarios.
The three methods that use both new data and synthetic kinematic trajectories to supervise both $a_{f,t}$ and $a_{m,t}$ (B2, B3, B4) show a relatively steady decrease in average task completion times. However, compared to ILSA, they improve performance at a slower rate, indicating that they are not utilizing new data as efficiently as ILSA.
(a) Cereal pouring task
(b) Pill bottle storage task
Figure 5: Task completion times averaged over all participants in each group. Error bars show standard deviations.
IV-C User Study
We conducted a user study subject to a university-approved IRB protocol with 20 participants, none of whom reported having prior experience teleoperating robots. We test the following two hypotheses:
  • H1: ILSA supports faster robot manipulation task completion times and easier manipulator control as compared to pure teleoperation.
  • H2: ILSA improves task performance over time as a user repeatedly performs a task, in contrast to a static shared autonomy method.
Participants were divided into two groups. Group 1, consisting of 10 participants, used ILSA and pure teleoperation, while Group 2, consisting of 10 participants, used teleoperation and a static shared autonomy policy, which was pretrained on the initial synthetic kinematic trajectories and kept static across the four task trials.
Participants completed the task trials twice with different methods in a counterbalanced order. While the within-subject comparison in Group 1 is sufficient to provide evidence towards H1, Group 2 is included to ensure that the reduction in task completion time with ILSA is not simply due to the user’s increasing familiarity with the shared autonomy system, which provides support for H2.
IV-C1 Task Completion Times
The task completion times of both groups using the varying methods are shown in Fig 5. For both tasks across all four trials, ILSA (dark blue bars) allowed for faster task completion in comparison to teleoperation (light blue bars). On average across all four trials, the cereal pouring task with ILSA took 10.2 seconds (12.97%) less than with teleoperation, while the pill bottle storage task with ILSA took 68.375 seconds (51.31%) less than with teleoperation.
A Wilcoxon signed-rank test on the combined task completion times from all four trials shows a statistically significant difference between ISLA and teleoperation ($p<0.001$ for both the cereal pouring and pill bottle storage tasks), providing evidence towards H1.
In addition, as shown in Fig 5, ILSA steadily reduced task completion times for both tasks. After only 3 trials of cereal pouring, ILSA decreased the average completion time from 92.1 seconds to 47.9 seconds, a 48.0% reduction. For the pill bottle storage task, the average time dropped from 82.0 seconds to 52.7 seconds, a 35.7% reduction, across all participants in Group 1.
To ensure that the reduction in task completion times with ILSA is not simply due to users’ increasing familiarity with the shared autonomy system, we compared results from group 1 with group 2. A Mann-Whitney U test on trial 4 shows no statistically significant difference between teleoperation in group 1 and group 2 ($p>0.05$ for both tasks). However, there is a statistically significant difference between ILSA and the static shared autonomy policy ($p<0.01$ for the cereal pouring task and $p<0.001$ for the pill bottle storage task), supporting H2.
IV-C2 User-reported Qualitative Metrics
In addition to capturing task completion time, we asked participants to report their perception of the robot control using 7-point Likert-type items (1 = strongly disagree; 4 = neutral; 7 = strongly agree) with items corresponding to “feel in control”, “low mental effort”, “intuitiveness”, “ease of task completion”, “smooth motions”, “precise motions”, and “no errors”. Participants’ median responses at trial 4 are shown in Fig. 6.
For both tasks, most participants reported a higher value on most metrics for ILSA compared to teleoperation. By applying a Wilcoxon signed-rank test, for the cereal pouring task, we observe a statistically significant difference between ILSA and teleoperation on “low mental effort” ($p<0.01$), “intuitiveness” ($p<0.05$), “ease of task completion” ($p<0.01$), “smooth motions” ($p<0.01$) and “no errors” ($p<0.05$). For the pill bottle storage task, we observe a statistically significant difference between ILSA and teleoperation on all 7 metrics ($p<0.05$ for “feel in control” and “precise motions”, $p<0.01$ for “low mental effort”, “intuitiveness”, “ease of task completion”, “smooth motions”, and “no errors”). These results support that users in general prefer ILSA over teleoperation, which provides evidence towards H1.
(a) Cereal pouring task
(b) Pill bottle storage task
Figure 6: Median Likert item responses for all methods after trial 4.
V CONCLUSIONS
In this paper, we introduce an innovative Incrementally Learned Shared Autonomy (ILSA) system that starts with pretraining on synthetic kinematic trajectories and then incrementally finetuned through user interactions. Our finetuning paradigm supports continuous improvement after each user interaction, balancing the integration of new data with the retention of prior knowledge. Through ablation and user studies, we compare ILSA to alternative designs, teleoperation, and a static shared autonomy method, demonstrating that ILSA is capable of enhancing task performance for complex multi-stage tasks over repeated task execution.